OpenStudy (anonymous):

a $60 jacket is discounted by 20%. when it doesn't sell, it is discounted another 20%. what is the new price?

zepdrix (zepdrix):

Hmm if it is discounted 20%, that means the new price should be 80% of the original right? If that doesn't make sense, let me know :D So let's multiply $60 by 80% to find the first discounted price. Converting to a decimal we get,\[\large 60 \times 0.8\]What value do you get for the first discounting? :O

OpenStudy (anonymous):

chow did you get 80%?

zepdrix (zepdrix):

The jacket originally costs 100% of it's price. If we SUBTRACT 20% from that price, we are left with 80% right? :D 100% - 20% = 80% You can do it the other way if you're more comfortable with it. \[\large 60 \times 0.2\]Then subtract that value from 60.

OpenStudy (anonymous):

ok, so 60 x .8 = $48 which is the first discounted price

zepdrix (zepdrix):

If the jacket doesn't sell, we discount it ANOTHER 20%. So if we take 80% of this already discounted jacket, it will effectively discount it another 20%.

zepdrix (zepdrix):

\[\large 48 \times 0.8\]

OpenStudy (anonymous):

= 38.40

zepdrix (zepdrix):

Yess good good! :) $38.40

zepdrix (zepdrix):

One of the things you might be asking yourself is, "Wait... if we took 20% off, and then 20% off again.. why didn't we just take 40% at the start?" Ahh good question my fishy friend! Because the 2nd time we took 20% off, we applied that discount to the previously discounted price. It's not the same as applying 20% and 20% to the original price.

OpenStudy (anonymous):

oh! i get it! taking 40% off would have given us an entirey different answer!

zepdrix (zepdrix):

Yesss :D heh

OpenStudy (anonymous):

thnkz so much!